Determine numerically the first derivative of function tabulated below,

at the point x=4

x: 1 2 4 8 10

y=f(x):0 1 5 21 27


1
Expert's answer
2021-04-20T02:16:54-0400



Newton,s dividend difference formula is


"f(x)=f(x_o)+(x-x_o)f(x_0,x_1)+(x-x_0)(x-x_1)f(x_0,x_1,x_2)+(x-x_2)(x-x_3)f(x_2,x_3)"


    "=f(1)_(x-1)\times 1+(x-1)(x-2)\times 3 +(x-4)(x-8)\times 12"

 

   "=x-1+3x^2-9x+6+12x^2-144x+384"


    "=15x^2-152x+389"


"f'(x)=30x-152\\



f'(4)=120-152=-32"